# OmniJet-: The first cross-task foundation model for particle physics

### v1

@@ Abstract
Foundation models are multi-dataset and multi-task machine learning methods that once pre-trained can be fine-tuned for a large variety of downstream applications. The successful development of such general-purpose models for physics data would be a major breakthrough as they could improve the achievable physics performance while at the same time drastically reduce the required amount of training time and data.
We report significant progress on this challenge on several fronts. First, a comprehensive set of evaluation methods is introduced to judge the quality of an encoding from physics data into a representation suitable for the autoregressive generation of particle jets with transformer architectures (the common backbone of foundation models). These measures motivate the choice of a higher-fidelity tokenization compared to previous works. Finally, we demonstrate transfer learning between an unsupervised problem (jet generation) and a classic supervised task (jet tagging) with our new OmniJet- model. This is the first successful transfer between two different and actively studied classes of tasks and constitutes a major step in the building of foundation models for particle physics.
@@ I Introduction
Beyond improving the physics performance by e.g. increasing selection accuracy of classification tasks, foundation models also address the other major issue currently facing particle physics: limited computing resources in the face of an ever increasing amount of dataĀ Albrecht etĀ al. (2019); Boehnlein etĀ al. (2022). This problem has already spawned the development of e.g. increasingly high fidelity models for the simulation of calorimeters, as well as techniques for the speeding up of other Monte Carlo simulationsĀ Paganini etĀ al. (2018); Buhmann etĀ al. (2021, 2023a); Adelmann etĀ al. (2022); Badger etĀ al. (2023); Hashemi and Krause (2023); Butter etĀ al. (2019); deĀ Oliveira etĀ al. (2017). By allowing the re-use of models across datasets and tasks, foundation models will play an important role in reducing this computational burden. Note that this effect will be further compounded by the potential for optimization in computing operations from one model used across multiple tasks.
Finally, using closely related architectures across different tasks inside one experiment, across experimental collaborations, and in the exchange with the theory community will make results easier to re-interpretĀ ara (2024); Bieringer etĀ al. (2024).
@@
Due to their flexibility demonstrated across language and other domains, transformers Vaswani etĀ al. (2017) are currently the most suitable candidate architecture for building foundation models for applications in particle physics. Taking inspiration from LLMs where sentences are generated autoregressively, recent efforts have demonstrated success with autoregressive generation of particle physics data, for example using the encoder part of the transformer to generate and jets Finke etĀ al. (2023) and using the decoder part of the transformer to generate +jets events Butter etĀ al. (2023). Ref. Heinrich etĀ al. (2024) demonstrated how a transformer backbone can be pre-trained using a BERT-like scheme where the model is trained to predict masked out jet constituents, resulting in an improvement of the performance when fine-tuning the backbone (with a new classification head) for jet tagging, especially at small training dataset sizes. Furthermore, Huang etĀ al. (2024) showed how a tokenized detector representation can be used in a BERT-like model for track reconstruction.
In this work, we will explore whether an autoregressive Generative Pretrained Transformer (GPT) model can be used as a foundation model for jet physics. However, the standard GPT constructions are not built to deal with continuous input data, but rather tokenized data. As point clouds are the most versatile representation of physics data Komiske etĀ al. (2019); Buhmann etĀ al. (2023b); Kasieczka etĀ al. (2019a); Buhmann etĀ al. (2023c, a) and can incorporate both event level information, jet substructure, and even low-level detector signals, finding a suitable input transformation for point clouds to tokens is the most pressing problem. Various tokenization strategies have been explored, for example using a simple mapping based on binning the input space in Finke etĀ al. (2023), a Gaussian mixture model in Butter etĀ al. (2023), and using an additional conditional embedding network in Heinrich etĀ al. (2024).
@@
Using this representation, we will first demonstrate training a generative model for jets as tokens in an unsupervised way for the JetClass Qu etĀ al. (2022b) dataset. Compared to Finke etĀ al. (2023), the core of our architecture is a transformer-decoder, not a transformer-encoder.
@@ II.1 Dataset
Jets are selected if they have a transverse momentum of and a pseudorapidity of . Additionally, truth-level matching is performed for all classes except and only jets that contain all the decay products of the boson or top quark are included. The resulting dataset contains 100M jets for training, 5M jets for validation, and 20M jets for testing.
In this work, only the kinematic information per particle (, , ) is used while the particle mass is approximated as zero. Next, the azimuth angle and the pseudorapidity are pre-processed to be relative to the jet axis111The difference in is signed and rectified to through . We handle those calculations using the scikit-hep/vectorĀ Schreiner etĀ al. (2023) and scikit-hep/awkwardĀ Pivarski etĀ al. (2024) libraries. :
Finally, we apply the cuts and to remove a very small fraction of low-energy constituents at the periphery and use up to 128 particles per jet.
@@ III.1 Token quality
Next, we consider distributions at the jet level to judge the quality of the tokenization. For this and the following studies, jets are mapped into token space, and then mapped back to physical space to quantify the loss in information. FigureĀ 4Ā (left) shows the jet mass combined for all classes in the dataset, as was done inĀ Heinrich etĀ al. (2024). As observed there, the conditional tokenization with 512 tokens already shows a good agreement between initial and reconstructed mass, with worse performance for the unconditional tokenization.
@@
However, this inclusive distribution might hide differences at the level of individual classes and jets. We therefore also consider the difference in mass for jets before and after tokenization and reconstruction FigureĀ 4Ā (center) for jets. The unconditional tokenization leads to a systematic shift of approximately 15Ā GeV, while the conditional tokenization is well centered at zero. Increasing the codebook size from 512 to 8192 tokens substantially improves the resolution. This behavior is even more pronounced when considering the -subjettinessĀ Thaler and VanĀ Tilburg (2011) ratio in FigureĀ 4Ā (right). Both conditional and unconditional tokenization with 512 tokens results in shifted distributions, while the larger codebook size of 8192 recovers a peak close to zero. Furthermore, while the mass resolution of the conditional tokens is already centered close to zero when using a codebook size of 512, the width of the distribution improves drastically from Ā GeV to Ā GeV when moving from a codebook size of 512 to a codebook size of 8192, where corresponds to the standard deviation obtained from fitting a normal distribution to the mass resolution histograms. A similar behavior can be observed for other classes, where in some cases, depending on the jet observable and the jet type, the effect is even more extreme. Finally, the binning approach with a 21x21x21 linear bins in the input features of our VQ-VAE comes close to the mass resolution of the conditional tokens with a codebook size of 8192, while the resolution of the subjettiness ratio is notably worse. Moreover, while this binning approach with 9162 tokens leads to reasonable resolution of the jets shown in FigureĀ 4, we found quite drastic mismodeling for and jets with such small codebook sizes222As expected, the resolution of the binning approach automatically leads to good resolution when the codebook size (i.e. the the number of bins) is increased to a sufficiently large number. We found that around tokens (corresponding to a 40x40x40 grid) offer similar resolution as conditional tokenization with a codebook size of 8192. . The distributions and the corresponding resolutions of the jet mass, jet , as well as the subjettiness ratios and are shown for all ten jet types individually in AppendixĀ B. Overall, the highest fidelity is achieved by conditional tokenization with a marked improvement from increasing the codebook size from 512 to 8192.
@@
Finally, we quantify the information loss that comes with the tokenization by training multi-class classifiers to distinguish between the ten jet types present in the dataset. The classifiers are trained once with the original inputs, and once with the inputs after undergoing tokenization and subsequent reconstruction back into physical space. This procedure allows a direct comparison of how the loss in resolution affects reconstruction performance. We utilize two standard classifier architectures: DeepSetsĀ Zaheer etĀ al. (2018); Komiske etĀ al. (2019) (i.e. without particle interactions) and TransformerĀ Vaswani etĀ al. (2017); Shleifer etĀ al. (2021) (i.e. with particle interactions) and perform this study for four different codebook sizes from 512 to 8192 tokens for the conditional tokenization approach. Details about the classifier trainings can be found in subsectionĀ A.2. Note that this approach is similar in spirit to the classifier metric proposed inĀ Krause and Shih (2023); Das etĀ al. (2023) but tests a multi-class classifier trained on these samples individually, as opposed to judging how well a classifier might distinguish original and reconstructed samples. This is necessary as e.g. points at fixed positions would be distinguishable from the original with close to perfect accuracy, rendering the test less useful.
@@ III.2 Jet generation
After training the transformer backbone with the generative head, it can be used for autoregressive generation as described in subsectionĀ II.3. The model was trained on three separate datasets: only, only, and and combined. This section will describe the combined version, since this is the model that is used for transfer learning. For a discussion of single-jet type generative results, including a comparison to the EPiC-FM method ofĀ Birk etĀ al. (2023), see appendix AppendixĀ C. 48Ā 000 events were generated from the combined model. These events contain tokens, which are then decoded back to physical space using the VQ-VAE decoder.
@@ III.3 Transfer learning from generation to classification
To evaluate the ability of the model to generalize from generating jets to classifying them, we focus on the task of hadronic top quark taggingĀ Kasieczka etĀ al. (2019a), i.e. distinguishing and jets on the JetClassĀ Qu etĀ al. (2022b) dataset. For this test, the Next-token prediction head is replaced by a Classification head while the transformer backbone is retained. We compare three training strategies: training the full architecture with randomly initialized weights (termed from scratch) which does not use transfer-learning and corresponds to the baseline, and two versions of fine-tuning the model obtained in the generative training step. In the Fine-tuning run, both the pre-trained backbone weights and the randomly initialized classification task head weights are allowed to float in the training, while in Fine-tuning (backbone fixed) only the classification task head is allowed to change.
The results of these training runs are presented in FigureĀ 7 as a function of the number of training examples provided to the model. We observe a significant gain in classification accuracy of both fine-tuning approaches compared to the baseline, leading to up to 15 percentage-points higher accuracy for small number of training jets, and outperforming by a few percentage-points at the highest training sample size. The difference between the two fine-tuning strategies is relatively small, with the more open training performing slightly better. Put differently, the generative pre-trained model achieves the same accuracy around 84% with 100 training examples for which the model that is trained from scratch requires 10ā000 examples.
@@ IV Conclusion
Most importantly, we report the generalization capability of OmniJet- from learning to generate jets in an unsupervised way, to the supervised classification between and jets. Overall, the fine-tuned model outperformed training from scratch for all values of training examples, often by a significant margin. For example, for 1000 training jets, the fine-tuned model achieves an accuracy of approximately 90%, compared to around 74% for the freshly initialized model. While the two approaches seem to converge, even at the highest training size of 2 million jets the fine-tuned approach maintains a lead of a few percentage points. Finally, it even provides a non-trivial classification accuracy of 84%, even when trained only on 100 jets, emphasizing the value of foundation models for problems with few available labeled training examples. While other types of transfer have been demonstrated previously, this is the first time that the unification across the two big classes of tasks ā classification and generation ā has been achieved.
@@ A.4 Transfer learning
To perform the transfer learning from the generative task to the classification task, we change the head of the OmniJet- model to the classification head and load the weights of the backbone trained for the generative task. We explore two variations of fine-tuning the pre-trained backbone to the classification task: training all weights of the model with the same learning rate (referred to as Fine-tuning in subsectionĀ III.3) and keeping the weights of the backbone fixed at the state obtained from the generative model (referred to as Fine-tuning (backbone fixed) in subsectionĀ III.3). For the From scratch trainings we start the training with randomly initialized weights of the whole model. The training is performed with the AdamWĀ Loshchilov and Hutter (2019) optimizer with a constant learning rate of 0.0001 and weight decay 0.01. Since those trainings, depending on the size of the training dataset, tend to show overfitting quite quickly, we stop those trainings when the validation loss does not improve for multiple epochs. The threshold of this early stopping is adjusted to the training dataset size, with a patience of 20 epochs for a training dataset of , and jets, a patience of 10 epochs for a training dataset size of and , and a patience of 5 for trainings with training jets. For each training dataset size we run 5 trainings with different random seeds and the epoch with the smallest validation loss is chosen for evaluation.
@@ Appendix B Token quality
Additional plots of the jet mass, the jet transverse momentum, as well as the subjettiness ratios are shown in FigureĀ 8, FigureĀ 9, FigureĀ 11 and FigureĀ 10.

### v2

@@ Abstract
Foundation models are multi-dataset and multi-task machine learning methods that once pre-trained can be fine-tuned for a large variety of downstream applications. The successful development of such general-purpose models for physics data would be a major breakthrough as they could improve the achievable physics performance while at the same time drastically reduce the required amount of training time and data. We report significant progress on this challenge on several fronts. First, a comprehensive set of evaluation methods is introduced to judge the quality of an encoding from physics data into a representation suitable for the autoregressive generation of particle jets with transformer architectures (the common backbone of foundation models). These measures motivate the choice of a higher-fidelity tokenization compared to previous works. Finally, we demonstrate transfer learning between an unsupervised problem (jet generation) and a classic supervised task (jet tagging) with our new OmniJet- model. This is the first successful transfer between two different and actively studied classes of tasks and constitutes a major step in the building of foundation models for particle physics.
@@ I Introduction
Beyond improving the physics performance by e.g. increasing selection accuracy of classification tasks, foundation models also address the other major issue currently facing particle physics: limited computing resources in the face of an ever increasing amount of dataĀ Albrecht etĀ al. (2019); Boehnlein etĀ al. (2022). This problem has already spawned the development of e.g. increasingly high fidelity models for the simulation of calorimeters, as well as techniques for the speeding up of other Monte Carlo simulationsĀ Paganini etĀ al. (2018); Buhmann etĀ al. (2021, 2024); Adelmann etĀ al. (2022); Badger etĀ al. (2023); Hashemi and Krause (2024); Butter etĀ al. (2019); deĀ Oliveira etĀ al. (2017). By allowing the re-use of models across datasets and tasks, foundation models will play an important role in reducing this computational burden. Note that this effect will be further compounded by the potential for optimization in computing operations from one model used across multiple tasks.
Finally, using closely related architectures across different tasks inside one experiment, across experimental collaborations, and in the exchange with the theory community will make results easier to re-interpretĀ Araz etĀ al. (2024); Bieringer etĀ al. (2024).
@@
Due to their flexibility demonstrated across language and other domains, transformers Vaswani etĀ al. (2017) are currently the most suitable candidate architecture for building foundation models for applications in particle physics. Taking inspiration from LLMs where sentences are generated autoregressively, recent efforts have demonstrated success with autoregressive generation of particle physics data, for example using a transformer to generate and jets Finke etĀ al. (2023) or to generate +jets events Butter etĀ al. (2023). In Heinrich etĀ al. (2024) it was demonstrated how a transformer backbone can be pre-trained using a BERT-like scheme where the model is trained to predict masked out jet constituents, resulting in an improvement of the performance when fine-tuning the backbone (with a new classification head) for jet tagging, especially at small training dataset sizes. Furthermore, Huang etĀ al. (2024) showed how a tokenized detector representation can be used in a BERT-like model for track reconstruction, and Hashemi etĀ al. (2024) used tokenization of detector images together with a transformer to capture the context within a collider event for subsequent generation111In addition, shortly after this paper was out, Harris etĀ al. (2024) demonstrated contrastive learning with re-simulated events as a suitable pre-training task..
In this work, we will explore whether an autoregressive Generative Pretrained Transformer (GPT) model can be used as a foundation model for jet physics. However, the standard GPT constructions are not built to deal with continuous input data, but rather tokenized data. As point clouds are the most versatile representation of physics data Komiske etĀ al. (2019); Buhmann etĀ al. (2023a); Kasieczka etĀ al. (2019a); Buhmann etĀ al. (2023b, 2024) and can incorporate both event level information, jet substructure, and even low-level detector signals, finding a suitable input transformation for point clouds to tokens is the most pressing problem. Various tokenization strategies have been explored, for example using a simple mapping based on binning the input space in Finke etĀ al. (2023), a Gaussian mixture model in Butter etĀ al. (2023), and using an additional conditional embedding network in Heinrich etĀ al. (2024).
@@
Using this representation, we will first demonstrate training a generative model for jets as tokens in an unsupervised way for the JetClass Qu etĀ al. (2022b) dataset.
@@ II.1 Dataset
Jets are selected if they have a transverse momentum of and a pseudorapidity of . Additionally, truth-level matching is performed for all classes except and only jets that contain all the decay products of the boson or top quark are included. The resulting dataset contains 100M jets for training, 5M jets for validation, and 20M jets for testing, with equal numbers of jets for each class. This split into training, validation, and test sets corresponds to the default split used in the JetClass dataset, in order to allow for a straightforward comparison with previous and future works.
In this work, only the kinematic information per particle (, , ) is used while the particle mass is approximated as zero. Next, the azimuth angle and the pseudorapidity are pre-processed to be relative to the jet axis222The difference in is signed and rectified to through . We handle those calculations using the scikit-hep/vectorĀ Schreiner etĀ al. (2023) and scikit-hep/awkwardĀ Pivarski etĀ al. (2024) packages. :
Finally, we apply the cuts and to remove a very small fraction of low-energy constituents at the periphery and use up to 128 particles per jet. The jet constituents are ordered by in descending order.
@@ III.1 Token quality
Next, we consider distributions at the jet level to judge the quality of the tokenization. For this and the following studies, jets are mapped into token space, and then mapped back to physical space to quantify the loss in information. FigureĀ 4Ā (left) shows the jet mass combined for all classes in the dataset, as was done inĀ Heinrich etĀ al. (2024). We observe a slight slope in the ratio plot for the unconditional tokens with codebook size 512, which is due to the underestimation of the masses (see also fig 8). The other approaches have a ratio close to 1 across the bulk of the histogram, and only show deviations at the tails where the statistics is low.
@@
However, this inclusive distribution might hide differences at the level of individual classes and jets. We therefore also consider the difference in mass for jets before and after tokenization and reconstruction FigureĀ 4Ā (center) for jets. The unconditional tokenization leads to a systematic shift of approximately 15Ā GeV, while the conditional tokenization is well centered at zero. Increasing the codebook size from 512 to 8192 tokens substantially improves the resolution. This behavior is even more pronounced when considering the -subjettinessĀ Thaler and VanĀ Tilburg (2011) ratio in FigureĀ 4Ā (right). Both conditional and unconditional tokenization with 512 tokens results in drastically shifted distributions, while the larger codebook size of 8192 recovers a peak close to zero for the mass difference. Furthermore, using 8192 tokens moves the peak of the difference closer to zero, while a small bias towards smaller in reconstructed jets is still present. In addition to that, while the mass resolution of the conditional tokens is already centered close to zero when using a codebook size of 512, the width of the distribution improves drastically from Ā GeV to Ā GeV when moving from a codebook size of 512 to a codebook size of 8192, where corresponds to the standard deviation obtained from fitting a normal distribution to the mass resolution histograms. A similar behavior can be observed for other classes, where in some cases, depending on the jet observable and the jet type, the effect is even more extreme. Finally, the binning approach with a 21x21x21 linear bins333The binning approach with 9261 tokens in a 21x21x21 grid is used as a reference, as it is the closest (but larger) number of bins (obtained with a NxNxN grid) compared to the codebook size of 8192 tokens. in the input features of our VQ-VAE comes close to the mass resolution of the conditional tokens with a codebook size of 8192, while the resolution of the subjettiness ratio is notably worse. Moreover, while this binning approach with 9162 tokens leads to reasonable resolution of the jets shown in FigureĀ 4, we found quite drastic mismodeling for and jets with such small codebook sizes444As expected, the resolution of the binning approach automatically leads to good resolution when the codebook size (i.e. the number of bins) is increased to a sufficiently large number. We found that around tokens (corresponding to a 40x40x40 grid) offer similar resolution as conditional tokenization with a codebook size of 8192. . The distributions and the corresponding resolutions of the jet mass, jet , as well as the subjettiness ratios and are shown for all ten jet types individually in AppendixĀ B. Overall, the highest fidelity is achieved by conditional tokenization with a marked improvement from increasing the codebook size from 512 to 8192.
@@
Finally, we quantify the information loss that comes with the tokenization by training multi-class classifiers to distinguish between the ten jet types present in the dataset. The classifiers are trained once with the original inputs, and once with the inputs after undergoing tokenization and subsequent reconstruction back into physical space. This procedure allows a direct comparison of how the loss in resolution affects reconstruction performance. We utilize two standard classifier architectures: DeepĀ SetsĀ Zaheer etĀ al. (2018); Komiske etĀ al. (2019) (i.e. without particle interactions) and TransformerĀ Vaswani etĀ al. (2017); Shleifer etĀ al. (2021) (i.e. with particle interactions) and perform this study for four different codebook sizes from 512 to 8192 tokens for the conditional tokenization approach. Details about the classifier trainings can be found in subsectionĀ A.2. Note that this approach is similar in spirit to the classifier metric proposed inĀ Krause and Shih (2023); Das etĀ al. (2023) but tests a multi-class classifier trained on these samples individually, as opposed to judging how well a classifier might distinguish original and reconstructed samples. This is necessary as e.g. points at fixed positions would be distinguishable from the original with close to perfect accuracy, rendering the test less useful.
@@ III.2 Jet generation
After training the transformer backbone with the generative head, it can be used for autoregressive generation as described in subsectionĀ II.3. The model was trained on three separate datasets: only, only, and and combined. This section will describe the combined version, since this is the model that is used for transfer learning. For a discussion of single-jet type generative results, including a comparison to the EPiC-FM model ofĀ Birk etĀ al. (2023), see appendix AppendixĀ C. 48ā000 events were generated from the combined model. These events contain tokens, which are then decoded back to physical space using the VQ-VAE decoder.
@@ III.3 Transfer learning from generation to classification
To evaluate the ability of the model to generalize from generating jets to classifying them, we focus on the task of hadronic top quark taggingĀ Kasieczka etĀ al. (2017, 2019a), i.e. distinguishing and jets on the JetClassĀ Qu etĀ al. (2022b) dataset. For this test, the Next-token prediction head is replaced by a Classification head while the transformer backbone is retained. We compare three training strategies: training the full architecture with randomly initialized weights (termed from scratch) which does not use transfer-learning and corresponds to the baseline, and two versions of fine-tuning the model obtained in the generative training step. In the Fine-tuning run, both the pre-trained backbone weights and the randomly initialized classification task head weights are allowed to float in the training, while in Fine-tuning (backbone fixed) only the classification task head is allowed to change.
The results of these training runs are presented in FigureĀ 7 as a function of the number of training examples provided to the model. We observe a significant gain in classification accuracy of both fine-tuning approaches compared to the baseline, leading to up to 17 percentage-points higher accuracy for small number of training jets, and outperforming by a few percentage-points at the highest training sample size. The difference between the two fine-tuning strategies is relatively small, with the more open training performing slightly better. Put differently, the generative pre-trained model achieves an accuracy of around 85% with 100 training examples for which the model that is trained from scratch requires more than 10ā000 examples.
@@ IV Conclusion
Most importantly, we report the generalization capability of OmniJet- from learning to generate jets in an unsupervised way, to the supervised classification between and jets. Overall, the fine-tuned model outperformed training from scratch for all values of training examples, often by a significant margin. For example, for 1000 training jets, the fine-tuned model achieves an accuracy of approximately 90%, compared to around 75% for the freshly initialized model. While the two approaches seem to converge, even at the highest training size of 2 million jets the fine-tuned approach maintains a lead of a few percentage points. Finally, it even provides a non-trivial classification accuracy of 85%, even when trained only on 100 jets, emphasizing the value of foundation models for problems with few available labeled training examples. Since the backbone trains completely without labels, it allows us to use a large unlabeled dataset to boost the performance of a smaller labeled one. While other types of transfer have been demonstrated previously, this is the first time that the unification across the two big classes of tasks ā classification and generation ā has been achieved.
@@
Code availability
The code for OmniJet- is publicly available at https://github.com/uhh-pd-ml/omnijet_alpha.
@@ Appendix A Model details and hyperparameters
Different hyperparameter configurations were tested for the individual model components of OmniJet-. The configurations presented in the following were found to lead to stable results. However, no extensive hyperparameter optimization was performed.
@@ A.4 Transfer learning
To perform the transfer learning from the generative task to the classification task, we change the head of the OmniJet- model to the classification head and load the weights of the backbone trained for the generative task. We explore two variations of fine-tuning the pre-trained backbone to the classification task: training all weights of the model with the same learning rate (referred to as Fine-tuning in subsectionĀ III.3) and keeping the weights of the backbone fixed at the state obtained from the generative model (referred to as Fine-tuning (backbone fixed) in subsectionĀ III.3). For the From scratch trainings we start the training with randomly initialized weights of the whole model. The training is performed with the AdamWĀ Loshchilov and Hutter (2019) optimizer with a constant learning rate of 0.0001 and weight decay 0.01. The transfer learning studies are performed on the test dataset from the default train-val-test split of the JetClass dataset, which includes 2M jets per jet-type, leading to 4M jets in total for our studies with and jets. We split those 4M jets into 2M jets for training, and 1M jets for validation and testing, respectively. Since those trainings, depending on the size of the training dataset, tend to show overfitting quite quickly, we stop those trainings when the validation loss does not improve for multiple epochs. The threshold of this early stopping is adjusted to the training dataset size, with a patience of 20 epochs for a training dataset of , and jets, a patience of 10 epochs for a training dataset size of and , and a patience of 5 for trainings with training jets. For each training dataset size we run 5 trainings with different random seeds and the epoch with the smallest validation loss is chosen for evaluation.
@@ Appendix B Token quality
Additional plots of the jet mass, the jet transverse momentum, as well as the subjettiness ratios are shown in FigureĀ 8, FigureĀ 9, FigureĀ 10, FigureĀ 11, FigureĀ 12, FigureĀ 13, FigureĀ 14 and FigureĀ 15.